### v1

# Activation thresholds and expressiveness of polynomial neural networks

Abstract
Polynomial neural networks have been implemented in a range of applications and present an advantageous framework for theoretical machine learning. A polynomial neural network of fixed architecture and activation degree gives an algebraic map from the network’s weights to a set of polynomials. The image of this map is the space of functions representable by the network. Its Zariski closure is an affine variety known as a neurovariety. The dimension of a polynomial neural network’s neurovariety provides a measure of its expressivity. In this work, we introduce the notion of the activation threshold of a network architecture which expresses when the dimension of a neurovariety achieves its theoretical maximum. In addition, we prove expressiveness results for polynomial neural networks with equi-width architectures.
Introduction
Polynomial neural networks are important in applications and theoretical machine learning. The function spaces and dimensions of neurovarieties for deep linear networks have been studied, and new developments in the polynomial neural network setting have appeared. In particular, results on the choice of the activation degree and the dimension of the neurovariety have improved our understanding of the optimization process of these neural networks and the ability of shallow and deep neural networks to replicate target functions [21, 27].
These theoretical results possess relevant implications. For appropriate datasets, polynomial activation functions can reduce model complexity and computational costs by introducing higher-order interactions between inputs, making it possible to model non-linear phenomena more efficiently. Moreover, polynomial neural networks have been found to perform well in practice in high-impact fields such as healthcare and finance. Polynomial neural networks have been successfully used to predict epidemic behavior [13], quantify the natural frequency of materials [11], analyze financial time series data [14, 25], and improve three-dimensional shape representation for computer vision [32].
The expressiveness of artificial neural networks of different structures and activation has been studied from several perspectives [23, 24, 29]. Substantial literature has shown the expressive power of deep networks, from the perspective of number of linear regions [24], universal approximation and VC dimensions [31, 29], a measure of complexity called “trajectory length” [23], and the exact class of functions representable by ReLU networks of various depths [18]. In another direction, [27] studied the effect of width of convolutional neural networks on the loss landscape, and [22, 17] provided bounds on network width for universal approximation for ReLU networks, while [19] provided such a bound for arbitrary activation functions. The expressiveness of neural networks has also been studied via their topological complexity, for example [4, 16]. In particular, [28] showed that networks with non-increasing architecture (each layer is no wider than the previous) result in a connected decision region. ReLU expressivity results have also been studied using algebraic geometry, specifically tropical geometry [6, 16, 18].
An intuitive way of characterizing the expressiveness of a neural network is by the dimension of the space of representable functions as one allows its parameters to vary. For sigmoid activation functions this has been done in [1], for ReLU activation and non-increasing architectures (as well as a few others) this was done by [7, 16]. For polynomial activation functions, [20, 21] studied this quantity and raised various conjectures, one of which we will resolve in the current paper.
This paper is structured as follows. In Section 1, we review key definitions for polynomial neural networks and introduce the notion of activation thresholds (8). Section 2 determines the expressivity of polynomial neural networks with high activation degree and proves the existence of the activation threshold. The main result resolves a conjecture by Kileel, Trager, and Bruna [20, Conjecture 16] and Kubjas, Li, Wiesmann [21, Conjecture 5.2]. In Section 3, we prove the expressivity of networks with equi-width architecture by computing the activation threshold— in particular, we remove any hypothesis involving “sufficiently high learning degree” by showing the activation threshold of these networks is one. In Section 4 we provide an outlook for the future.
1 Polynomial neural networks and neurovarieties
For consistency, the notation introduced in this section is intended to agree with [20] and [21] as much as possible.
An $L$-layer feedforward neural network $F_{\boldsymbol{\theta}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ is a composition of affine-linear maps $f_{i}:\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}}$ and non-linear maps $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$,
$F_{\boldsymbol{\theta}}(x)=(f_{L}\circ\sigma_{L-1}\circ f_{L-1}\circ\cdots\circ f_{2}\circ\sigma_{1}\circ f_{1})(x).$
The architecture of the neural network $F_{\boldsymbol{\theta}}$ is the sequence $\mathbf{d}=(d_{0},\dotsc,d_{L})$. Here, the affine-linear maps are given by
$f_{i}(x):\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}},\quad x\mapsto W_{i}x+b_{i},$
where the weights of the neural network are the matrices $W_{1},\dots,W_{L}$ and the biases are the vectors $b_{1},\dotsc,b_{L}$. The parameter set ${\boldsymbol{\theta}}$ consists of these weight matrices and biases. The activation map $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ is given coordinate-wise by the activation function.
Deep learning consists of approximating a target function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by a neural network $F_{\boldsymbol{\theta}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ of a chosen architecture $\mathbf{d}=(d_{0},\dotsc,d_{L})$ and activation function. That is, deep learning concerns recovering parameters ${\boldsymbol{\theta}}$ for which $F_{\boldsymbol{\theta}}$ most closely resembles $F$. The map $F_{\boldsymbol{\theta}}$ is associated to the tuple of parameters ${\boldsymbol{\theta}}$ by the parameter map
$\Psi_{\mathbf{d},\sigma}:\mathbb{R}^{N}\to\text{Fun}(\mathbb{R}^{d_{0}},\mathbb{R}^{d_{L}}),\quad{\boldsymbol{\theta}}\mapsto F_{\boldsymbol{\theta}}.$ (1)
In this article, we are interested when the activation function consists of polynomials. The most important case are pure powers, and this leads to the following definition.
Definition 1 (Polynomial neural network).
A polynomial neural network $p_{\textbf{w}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ with fixed activation degree $r$ and architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ is a feedforward neural network of the form
$p_{\mathbf{w}}(x)=(W_{L}\circ\sigma_{L-1}\circ W_{L-1}\circ\sigma_{L-2}\circ\cdots\circ\sigma_{1}\circ W_{1})(x)$
where $W_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and the activation maps $\sigma_{i}(x):\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ are given by coordinate-wise exponentiation to the $r$-th power,
$\sigma_{i}(x):=(x_{1}^{r},\dots,x_{d_{i}}^{r}).$
The parameters $\mathbf{w}$ are the entries in the matrices $W_{i}$, so that
$\mathbf{w}=(W_{1},W_{2},\dots,W_{L}).$
Remark 2.
In the definition of polynomial neural network, the biases are taken to be zero. The affine-linear map $f_{i}$ in the $i$-th layer is then a linear map that we identify with the matrix $W_{i}$.
An $L$-layer polynomial neural network $p_{\mathbf{w}}$ with architecture $\mathbf{d}=(d_{0},\dotsc,d_{L})$ and activation degree $r$ is a tuple of homogeneous polynomials. The parameter map now takes the matrix tuple $(W_{1},W_{2},\dots,W_{L})$ to a tuple of degree $r^{L-1}$ homogeneous polynomials in $d_{0}$ variables of length $d_{L}$
$\Psi_{\mathbf{d},r}:\mathbb{R}^{d_{1}\times d_{0}}\times\cdots\times\mathbb{R}^{d_{L}\times d_{L-1}}\to(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}},\quad\mathbf{w}\mapsto p_{\mathbf{w}}=\begin{pmatrix}p_{\mathbf{w}_{1}}\\ \vdots\\ p_{\mathbf{w}_{d_{L}}}\end{pmatrix}.$ (2)
To specify an element in the image of $\Psi_{\mathbf{d},r}$, it suffices to identify its vector of coefficients in $\mathbb{R}^{d_{L}\binom{r^{L-1}+d_{0}-1}{d_{0}-1}}\simeq(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
Remark 3 (Special cases).
Note that if $d_{L}=1$, the co-domain of the parameter map is the space of degree $r^{L-1}$ homogeneous polynomials. If $L=1$ then the ambient space consists of linear maps from $\mathbb{R}^{d_{0}}$ to $\mathbb{R}^{d_{L}}$.
The neuromanifold or functional space associated to $p_{\mathbf{w}}$ is the family of all polynomial neural networks with fixed activation degree $r$ and architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ where the weights are allowed to vary.
Definition 4 (Neurovariety).
The image of $\Psi_{\mathbf{d},r}$ is the neuromanifold $\mathcal{M}_{\mathbf{d},r}$, which is a semialgebraic set in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$. The neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the Zariski closure of $\mathcal{M}_{\mathbf{d},r}$, and is an affine variety in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
The dimension of the neurovariety $\mathcal{V}_{\mathbf{d},r}$ provides a measure of the expressivity of a neural network because it quantifies the degrees of freedom of the space of functions the network produces [20]. In practice, this is the dimension of the neuromanifold $\mathcal{M}_{\mathbf{d},r}$; as $\mathcal{M}_{\mathbf{d},r}$ is semialgebraic, its dimension is the same as that of $\mathcal{V}_{\mathbf{d},r}$.
The size of the weight matrices is a primary factor in determining the dimension of the neurovariety. However, for all network architectures there exists a symmetry in the weight matrices, known as multi-homogeneity. It follows that the dimension of a fiber of the parameter map $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$. ([20], Lemma 13).
Lemma 5 ([20], Multi-homogeneity).
For all invertible diagonal matrices $D_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ and permutation matrices $P_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ ($i=1,\dots,L-1$), the parameter map $\Psi_{\mathbf{d},r}$ returns the same map under the replacement
$W_{1}$ $\leftarrow P_{1}D_{1}W_{1}$
$W_{2}$ $\leftarrow P_{2}D_{2}W_{2}D_{1}^{-r}P_{1}^{T}$
$\vdots$
$W_{L}$ $\leftarrow W_{L}D_{L-1}^{-r}P_{L-1}^{T}.$
where $T$ indicates matrix transpose. Thus the dimension of a generic pre-image of $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$.
Considering the difference of the size of the weight matrices and the dimension of the space of multi-homogeneities leads to the notion of the expected dimension of a neurovariety.
Definition 6 (Expected dimension).
The expected dimension of the neurovariety $\mathcal{V}_{\mathbf{d},r}$ is
$\text{edim}\mathcal{V}_{\mathbf{d},r}:=\min\left\{d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1}),\ d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}\right\}.$
This definition was introduced in [21] and in [20] it was shown that $\text{edim}\mathcal{V}_{\mathbf{d},r}$ is an upper bound on the dimension of the neurovariety. However, this expected dimension is not always equal to the dimension, as demonstrated by the following example.
Example 7.
The dimension of a neurovareity is subtle even in simple settings. For instance, when the architecture is $(d_{0},d_{1},1)$, a reinterpretation [20, Theorem 9] of the Alexander–Hirschowitz Theorem [2] determines which neurovarieties have the expected dimension. In [21], the authors coin the term defect for the difference between the dimension of a neurovariety from its expected dimension. Examples of $\mathcal{V}_{\mathbf{d},r}$ with defect are found in [21, Table 1] and in [20] as well.
Our main result proves a lower bound on the dimension of a neurovariety for sufficiently high activation degree by resolving a conjecture in [20]. We introduce the notion of the activation threshold, whose existence will follow from our main result Theorem 12.
Problem 8 (Activation threshold).
Given an architecture $\mathbf{d}$, does there exist a non-negative integer $\tilde{r}$ such that the following holds:
$\dim\mathcal{V}_{\mathbf{d},r}=\text{edim}\mathcal{V}_{\mathbf{d},r}\quad\text{for all}\quad r>\tilde{r}\,?$ (3)
Definition 9.
The activation threshold is the smallest $\tilde{r}$ such that (3) holds, if it exists. We denote the activation threshold of $\mathbf{d}$ by $\operatorname{ActThr}(\mathbf{d})$.
2 Expressivity for high activation degree
In this section we resolve the high activation degree conjecture of Kileel, Trager, and Bruna. Precisely, we prove [20, Conjecture 16] thereby completing the proof of [20, Theorem 14]. The key piece of our argument incorporates the number theoretic results by Newman and Slater in [26].
2.1 Powers of non-proportional polynomials
We generalize the result [26, Section 5, Equation (2)] of Newman-Slater to multivariate polynomials. The proof of this is via a simple reduction from multivariate polynomials to univariate polynomials.
Lemma 10 (Newman-Slater Corollary).
Let $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ denote multivariate polynomials that are pairwise non-proportional (for any $i\not=j$, there is no $\alpha\in\mathbb{C}$ such that $p_{i}=\alpha p_{j}$). If
$p_{1}^{r}+\dotsb+p_{k}^{r}=0,$
then $r\leq 8(k-1)^{2}-1$.
Proof.
Let $p_{1},\dotsc,p_{k}$ be given as above and fix a general line in $\mathbb{C}^{m}$ parameterized by $\ell(t)$. As $\sum_{i}p_{i}^{r}=0$, the restriction of the polynomials $p_{i}$ to this line yields the following equality of univariate polynomials $p_{1}(\ell(t)),\dotsc,p_{k}(\ell(t))$:
$(p_{1}(\ell(t)))^{r}+\dotsb+(p_{k}(\ell(t)))^{r}=0.$
As the $p_{i}$ are pairwise non-proportional and the line parameterized by $\ell(t)$ is general, the univariate polynomials $p_{i}(\ell(t))$ are also pairwise non-proportional as they share no common zeros. By the result of Newman and Slater [26, Section 5, Equation (2)], we conclude that $r\leq 8(k-1)^{2}-1$. ∎
We generalize Lemma 10 further—we demonstrate that given pairwise non-proportional multivariate polynomials $p_{1},\dotsc,p_{k}$, the polynomials $p_{1}^{r},\dotsc,p_{k}^{r}$ are linearly independent for all sufficiently large $r$.
Lemma 11.
Let $\mathbb{K}$ be a subfield of $\mathbb{C}$. Given integers $d,k$ there exists an integer $\tilde{r}=\tilde{r}(k)$ with the following property. If $r>\tilde{r}$ and $p_{1},\dots,p_{k}\in\mathbb{K}[x_{1},\dots,x_{d}]$ are pairwise non-proportional, then $p_{1}^{r},\dots,p_{k}^{r}$ are linearly independent (over $\mathbb{K}$). Moreover, $\tilde{r}=8(k-1)^{2}-1$ suffices.
Proof.
We remark that it suffices to consider the case that $\mathbb{K}=\mathbb{C}$. Indeed, if $p_{i}$ and $p_{j}$ are not linearly dependent over $\mathbb{K}$, then they are not linearly dependent over $\mathbb{C}$. For if there were a constant $\alpha\in\mathbb{C}$ such that $p_{i}-\alpha p_{j}=0$, then one finds that $\alpha\in\mathbb{K}$ by considering the coefficients of this difference. Further, linear independence of $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ over $\mathbb{C}$ implies linear independence over any subfield.
We prove the contrapositive statement when $\mathbb{K}=\mathbb{C}$. Given integers $d,k$, let $\tilde{r}=8(k-1)^{2}-1$ and $r\geq\tilde{r}$ be an integer. Fix pairwise non-proportional polynomials $p_{1},\dotsc,p_{k}\in\mathbb{C}[x_{1},\dotsc,x_{d}]$ with the property that the set $\{p_{1}^{r},\dotsc,p_{k}^{r}\}$ is linearly dependent over $\mathbb{C}$. Then there exists $\alpha_{1},\dots,\alpha_{k}\in\mathbb{C}$ such that
$\alpha_{1}p_{1}^{r}+\cdots+\alpha_{k}p_{k}^{r}=0.$
Let $\beta_{i}$ be a $r$th root of $\alpha_{i}$, so that
$(\beta_{1}p_{1})^{r}+\cdots+(\beta_{k}p_{k})^{r}=0.$
By Lemma 10, it follows that for some $i\neq j$, $\beta_{i}p_{i}$ and $\beta_{j}p_{j}$ are proportional. Thus, some $p_{i}$ and $p_{j}$ are proportional.
2.2 Deep networks and high activation degree
As noted in [20], Lemma 11 implies that for a fixed architecture $\mathbf{d}$, there exists $\tilde{r}$ such that for any activation degree $r>\tilde{r}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension [20, Theorem 14]. That is, Lemma 11 implies the existence of the activation threshold for any architecture $\mathbf{d}$. We recount the argument given in [20] for completeness.
Theorem 12.
For fixed $\mathbf{d}=(d_{0},\dotsc,d_{L})$ satisfying $d_{i}>1$ ($i=1,\dotsc,L$), the activation threshold $\operatorname{ActThr}(\mathbf{d})$ exists. In other words, there exists $\tilde{r}$ such that whenever $r>\tilde{r}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension,
$\dim\mathcal{V}_{\mathbf{d},r}=\text{edim}\mathcal{V}_{\mathbf{d},r}=d_{L}+\sum_{i=0}^{L-1}(d_{i}-1)d_{i+1}.$
Moreover, $\operatorname{ActThr}(\mathbf{d})\leq 8\cdot(2\cdot\max\{d_{1},\dots,d_{L-1}\}-1)^{2}-1$.
Proof.
Fix $r>\tilde{r}=8\cdot(2\cdot\max\{d_{1},\dotsc,d_{L-1}\}-1)^{2}-1$. We show by induction on the number of layers $L$ that for activation degree $r$, a general fiber of $\Psi_{\mathbf{d},r}$ is equal to the set of multi-homogeneities of any point of the fiber given by Lemma 5. The dimension of $\mathcal{V}_{\mathbf{d},r}$ is then equal to the expected dimension.
In the case that $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension as the map $\Psi_{\mathbf{d},r}$ is an isomorphism.
Now assume $L>1$ and fix general weights $W_{1},\dots,W_{L}$ and consider weights $\tilde{W}_{1},\dots,\tilde{W}_{L}$ such that
$(W_{L}\sigma_{L-1}W_{L-1}\dots\sigma_{1}W_{1})(x)=(\tilde{W}_{L}\sigma_{L-1}\tilde{W}_{L-1}\dots\sigma_{1}\tilde{W}_{1})(x).$ (4)
Denote the output of $(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x)$ by $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]$ and the output of $(\tilde{W}_{L-1}\sigma_{L-2}\tilde{W}_{L-2}\dotsb\tilde{W}_{2}\sigma_{1}\tilde{W}_{1})(x)$ by $[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$. Since the weights $W_{i}$ are general and $d_{i}>1$, the homogeneous polynomials $p_{\mathbf{w}_{i}}$ are pairwise non-proportional (there is a non-empty Zariski open set of the space of weight matrices where the outputs $p_{\mathbf{w}_{1}},\dotsc,p_{\mathbf{w}_{d_{L-1}}}$ are non-proportional). By examining the first coordinate of the outputs of (4) we obtain the decomposition
$w_{L_{11}}p_{\mathbf{w}_{1}}^{r}+\dots+w_{L_{1d_{L-1}}}p_{\mathbf{w}_{d_{L-1}}}^{r}=\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}_{\mathbf{w}_{d_{L-1}}}^{r}$ (5)
where $w_{L_{ij}}$ and $\tilde{w}_{L_{ij}}$ denote the $(i,j)$ entry of the weight matrices $W_{L}$ and $\tilde{W}_{L}$ respectively. Because we have taken $r>\tilde{r}$, Lemma 11 guarantees that (5) has two proportional summands. As $p_{\mathbf{w}_{i}}$ are not pairwise proportional, no two summands on the left side may be proportional. If two summands on the right side are proportional, we reduce the number of terms on the right side by combining these terms. Otherwise, by permuting terms as necessary, we may assume these proportional terms to occur as the first term on both sides. We may then scale so that $p_{\mathbf{w}_{1}}=\tilde{p}_{\mathbf{w}_{1}}$ and subtract $\tilde{w}_{L_{11}}\tilde{p}_{\mathbf{w}_{1}}^{r}$ to obtain
$(w_{L_{11}}-\tilde{w}_{L_{11}})p^{r}_{\mathbf{w}_{1}}+\dots+w_{L_{1d_{L-1}}}p^{r}_{\mathbf{w}_{d_{L-1}}}=\tilde{w}_{L_{12}}\tilde{p}_{\mathbf{w}_{2}}^{r}+\dots+\tilde{w}_{L_{1d_{L-1}}}\tilde{p}^{r}_{\mathbf{w}d_{L-1}}.$ (6)
We repeatedly apply Lemma 11 to iteratively reduce the number of terms on the right side, until the right side of (6) is zero. Then, the pairwise linear independence of the $p_{\mathbf{w}_{i}}$ and Lemma 11 imply that each coefficient is also zero. Hence, up to scaling and permutation, $[p_{\mathbf{w}_{1}},\dots,p_{\mathbf{w}_{d_{L-1}}}]=[\tilde{p}_{\mathbf{w}_{1}},\dots,\tilde{p}_{\mathbf{w}_{d_{L-1}}}]$ and the entries of $W_{L}$ and $\tilde{W}_{L}$ agree in the first row. This process is repeated for each row of $W_{L}$ and $\tilde{W}_{L}$ to show that they are equal (up to scaling and permutation). Therefore the fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(W_{1},\dots,W_{L}))$ has dimension $\sum_{i=1}^{L-1}d_{i}.$
3 Equi-width setting
The advantage of depth for neural networks is well-established through both empirical observations and theoretical results [24, 31, 23, 29]. Width also plays a significant role in network expressivity and improves the behavior of the loss surface by smoothing the optimization landscape of the loss function [18, 27, 28]. The study of networks with bounded width appears in the context of universal approximation theorems [17, 19, 22] as well.
For sufficiently high activation degree, a neurovariety has the expected dimension by Theorem 12. For specific architectures, better bounds on the activation threshold may be computed. In this section, we focus on determining the expressivity of polynomial neural networks with equi-width architecture, meaning the layers have equal widths, i.e., $d_{0}=d_{1}=\dots=d_{L}$ in Definition 1.
We begin with a proposition which will be utilized in the main result of this section.
Proposition 13.
Let $\mathbf{d}=(d_{0},\dotsc,d_{L})$ be an equi-width architecture such that $d:=d_{0}=d_{1}=\dotsb=d_{L}$ and let $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ be the activation map given by the coordinate-wise exponentiation to the $r$-th power. A polynomial neural network
$p_{\mathbf{w}}(x)=(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x)\in\mathcal{V}_{\mathbf{d},r}$
has only the trivial zero if and only if all $W_{i}$ are invertible.
Proof.
If each $W_{i}$ is invertible, then the preimage of the zero vector under each $W_{i}$ and $\sigma_{i}$ consists of only the zero vector. Thus, the polynomial neural network
$p_{\mathbf{w}}(x)=(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x)$
has only the trivial zero.
If some $W_{i}$ is singular, then let $i_{0}$ be the minimal index such that $W_{i_{0}}$ is singular and $v\in\ker W_{i_{0}}$ be a non-zero vector. For each $j<i_{0}$, the linear map $W_{j}$ is surjective, and each $\sigma_{j}$ is surjective as well. Thus, there exists non-zero $x_{\ast}\in\mathbb{C}^{d}$ such that
$(\sigma_{i_{0}-1}W_{i_{0}-1}\dotsb W_{2}\sigma_{1}W_{1})(x_{\ast})=v.$
As $v\in\ker W_{i_{0}}$, it follows that $p_{\mathbf{w}}(x_{\ast})=(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x_{\ast})=0$ and $x_{\ast}$ is a non-trivial zero. ∎
We now provide the main result of this section, that the activation threshold of an equi-width architecture is one. That is, for an equi-width architecture $\mathbf{d}$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r\geq 2$.
This is done by explicitly computing a single fiber of the parameter map $\Psi_{\mathbf{d},r}$. This dimension provides an upper bound on the generic dimension of a fiber, and so we obtain a lower bound on the dimension $\dim\mathcal{V}_{\mathbf{d},r}$. This lower bound is in fact equal to our upper bound given by the expected dimension, thereby proving the result.
Theorem 14.
If $\mathbf{d}=(d_{0},\dotsc,d_{L})$ is an equi-width architecture with $d=d_{0}=\dotsb=d_{L}$, $L>1$ and $d>1$, then the activation threshold is $\operatorname{ActThr}(\mathbf{d})=1$. That is, if $\mathbf{d}$ is equi-width with $L>1$ and $d>1$, then for all $r>1$ the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension
$\dim\mathcal{V}_{\mathbf{d},r}=Ld^{2}-(L-1)d.$
[…]
Proof.
[…]
We remark that when $L>1$, $\operatorname{ActThr}(\mathbf{d})>0$. Indeed, if $r=1$, then $\mathcal{V}_{\mathbf{d},r}$ is again the space of linear functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$ so that $\mathcal{V}_{\mathbf{d},r}=d^{2}$. Then $d>1$ implies that
$\dim\mathcal{V}_{\mathbf{d},r}=d^{2}<Ld^{2}-(L-1)d=\text{edim}\mathcal{V}_{\mathbf{d},r}.$
We show that when $L>1$ and $r>1$, a specific fiber of the parameter map
$\Psi_{\mathcal{V}_{\mathbf{d},r}}:\mathbb{R}^{d_{L}\times d_{L-1}}\times\dotsb\times\mathbb{R}^{d_{1}\times d_{0}}\to\mathcal{V}_{\mathbf{d},r}$
has dimension $(L-1)d$ and thus $\operatorname{ActThr}(\mathbf{d})=1$. Consider the neural network
$\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}=\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}).$
Precisely, we show that the fiber $\Psi_{\mathbf{d},r}^{-1}(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ is the set of multi-homogeneities given in Lemma 5.
To this end, let $W_{1},\dotsc,W_{L}$ be weight matrices such that
$(W_{L}\sigma_{L-1}W_{L-1}\dotsb W_{2}\sigma_{1}W_{1})(x)=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}.$
We will show by induction that the output of the $i$-th layer consists of scalar multiples of the monomials $x_{1}^{r^{i-1}},\dotsc,x_{d}^{r^{i-1}}$, which will imply that each $W_{i}$ is a (invertible) scaled permutation matrix. Consider the output of the $(L-1)$-th layer,
$\begin{pmatrix}p_{\mathbf{w}_{1}}(x)\\ \vdots\\ p_{\mathbf{w}_{d}}(x)\end{pmatrix}=(W_{L-1}\sigma_{L-2}W_{L-2}\dotsb W_{2}\sigma_{1}W_{1})(x).$
There is then an equality
$W_{L}\begin{pmatrix}p_{\mathbf{w}_{1}}^{r}(x)\\ \vdots\\ p_{\mathbf{w}_{d}}^{r}(x)\end{pmatrix}=\begin{pmatrix}x_{1}^{r^{L-1}}\\ \vdots\\ x_{d}^{r^{L-1}}\end{pmatrix}$ (7)
Since the polynomial neural network $(x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}})$ only has the trivial zero, Proposition 13 implies that $W_{L}$ is invertible. Multiplying both sides of (7) by $W_{L}^{-1}$, each $p_{\mathbf{w}_{i}}^{r}(x)$ is a linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$. For $r>1$, this is only possible if each $p_{\mathbf{w}_{1}}$ is a monomial. Indeed, no linear combination of $x_{1}^{r^{L-1}},\dotsc,x_{d}^{r^{L-1}}$ is a pure $r$-th power except for scalar multiples of these individual monomials. Thus each $p_{\mathbf{w}_{i}}^{r}$ is a scalar multiple of some $x_{j}^{r^{L-1}}$ and each $p_{\mathbf{w}_{i}}$ is a scalar multiple of some $x_{j}^{r^{L-2}}$.
It follows that $W_{L}$ is a scaled permutation matrix—it is an invertible matrix that sends a set of monomials to a set of scalar multiples of those monomials. As the output of the $(L-1)$-th layer is then a set of scalar multiples of $x_{i}^{r^{L-2}}$, we inductively apply this argument to show that the fiber has the form of Lemma 5.
The fiber $\Psi_{\mathbf{d},r}^{-1}(\Psi_{\mathbf{d},r}(I_{d},\dotsc,I_{d}))$ then consists of the set of multi-homogeneities and has the expected dimension $\sum_{j=1}^{L-1}d_{j}$, which is an upper bound on the generic dimension of a fiber. It follows that
$\text{edim}\mathcal{V}_{\mathbf{d},r}=\sum_{i=0}^{L-1}d_{i}d_{i+1}-\sum_{j=1}^{L-1}d_{j}$
is a lower bound for $\dim\mathcal{V}_{\mathbf{d},r}$. As $\text{edim}\mathcal{V}_{\mathbf{d},r}$ is also an upper bound, $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. ∎
4 Outlook
Bounding the activation threshold
Our bound on the activation threshold for polynomial neural networks is not universally optimal. We expect that a lemma of [15] used to prove Picard theorems for holomorphic maps into Fermat hypersurfaces will allow us to improve the bound in Theorem 12. Effective methods in algebraic geometry to compute the activation threshold of a given architecture also deserve to be studied. One possible approach is to formalize a connection to patterns of linear dependence among families of polynomials. Such patterns have notable examples from number theory independently due to A. H. Desboves [12] and Noam Elkies [10] and were studied in [30] for a finite set of arbitrary polynomials in terms of tickets. The ticket $\mathcal{T}(f)$ of a finite set of polynomials $F=\{f_{j}\}$ is defined as
$T(F)=\{m\in\mathbb{N}:\{f_{j}^{m}\}\text{ is linearly dependent}\}.$
By viewing the polynomial neural network recursively, layer by layer, it should be possible prove results about the expected dimension by translating tickets into the polynomial neural network setting.
Non-increasing architectures
One future direction in this research area is to determine the activation thresholds for non-increasing architectures $\mathbf{d}=(d_{0},\dotsc,d_{L})$ — those architectures satisfying $d_{0}\geq d_{1}\geq\dotsb\geq d_{L}$. By generalizing the proof of Theorem 14, one might hope to prove a conjecture of [21] that for a non-increasing architecture $\mathbf{d}$ with $d_{L}>1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. That is, prove that a non-increasing architecture $\mathbf{d}$ has activation threshold less than or equal to one.
Other networks
Another research direction is to introduce the notion of activation thresholds for other linear and non-linear neural networks. Permitting negative exponents in a polynomial activation map leads to the notion of rational neural networks [5]. There is experimental evidence that it is beneficial to choose rational activation maps that are very low degree. Defining activation thresholds for rational networks would provide a quantitative measure to make these observations rigorous. Rational neural networks are currently being investigated by Alexandros Grosdos et al. Certain linear networks may also admit the definition of an activation threshold as a means to study expressiveness. Tropical rational neural networks provide a lens to study networks with piecewise linear activation maps [3, 8, 9, 33]. Such networks encompass ReLU networks; a ReLU network of fixed architecture is described by a semialgebraic set inside the parameter space of tropical rational functions. In this context, tropical rational functions provide a means to define activation thresholds to frame and obtain results on expressiveness. Moreover, because the dimension of the neurovariety is preserved under tropicalization, activation thresholds may aid in translating expressiveness results for tropical neural networks to polynomial networks.